# Zhonghong Kuang Renmin U Yi Liu

### v1

@@ 5.1 Direct Training Cost
5.2 Comparison with Garicano and Rayo (2017)
In Garicano and Rayo (2017) (hereafter “GR”), the expert employs the novice to produce, but the novice initially possesses neither the productive knowledge nor the financial resources to acquire such knowledge. Because of that, the expert is intrinsically motivated to train the novice, but at the same time, he is concerned that the novice may (or may threaten to) stop producing for him and pursue an outside option whose value increases with the novice’s knowledge level. GR show that the expert-optimal contract is a multi-period apprenticeship, in which the novice is trained gradually over time and receives full knowledge after a finite number of periods. In the initial period, the expert offers a “knowledge gift,” despite the novice’s inability to produce; after that, the novice works without pay until he fully acquires the remaining knowledge; once the apprenticeship concludes, the novice starts to retain all the output from his production. In what follows, we highlight the key differences between our results and those in GR.
Finite vs. Perpetual Transfer Process
In GR, knowledge is fully transferred after a finite number of periods. This outcome arises because, as the amount of remaining knowledge diminishes, a threshold is eventually reached where the novice’s output in a single period exceeds the value of the remaining knowledge. At that point, the expert has an incentive to transfer all remaining knowledge at once to accelerate revenue collection. This result also hinges on the expert’s one-period commitment: If the expert can freely choose how much knowledge to transfer after the novice has produced, he would never transfer all remaining knowledge, knowing that the novice would then appropriate the entire surplus. In contrast, in our model with no commitment, the knowledge transfer process proceeds perpetually, with the novice’s knowledge level converging to a long-run limit. When players are somewhat impatient, a complete knowledge transfer is not even attainable at the time limit. [^20] This is true in our baseline model. In the model with retirement, knowledge is always fully transferred upon the expert’s retirement, regardless of the discount factor.
Knowledge Gift
Discount Factor and Speed of Knowledge Transfer
GR and our paper also yield distinct predictions regarding the effect of the discount factor on the speed of knowledge transfer. In GR, greater patience (i.e., a higher discount factor) leads to longer apprenticeships and slower knowledge transfer. This is because an increase in the discount factor raises the marginal value of knowledge to the novice, making him willing to work in exchange for smaller increments of knowledge per period. The expert exploits this by deliberately slowing the transfer of knowledge, thereby extending the apprentice phase during which he captures the full output. To the contrary, in our model, greater patience enhances the principal’s value from future knowledge transfers; as a result, the principal is willing to pay more to the expert in each period, which in turn supports larger and faster knowledge transfers. As discussed in Section 3.4, our prediction seems to align well with the comparison of long-run cooperation levels between the SAIC-GM joint venture and the SAIC-Volkswagen joint venture.
@@ 6 Concluding Remarks
We study how a principal should incentivize an expert to share specialized knowledge with a novice, when the expert is concerned about the future competition he has to face from a more capable novice. In the optimal relational contract between the principal and the expert, the expert is asked to transfer as much knowledge as he can to the novice, for free, in the initial period. The lack of commitment on both sides necessitates gradual and perpetual knowledge transfer later on, with the principal always compensating the expert for his future losses immediately after verifying the training he provided. When players are somewhat impatient, a complete knowledge transfer is not achievable even in the long run.
Our characterization matches well with the implementation details of the quid pro quo policy in the Chinese auto industry, a policy that has successfully facilitated significant technology transfers between various international companies and domestic automakers via joint ventures. The analysis can also be extended to a more complex environment that accounts for expert retirements and the career progression of novices.
While our model is stylized, we believe that the insights can shed light on a variety of applications, within and beyond the examples discussed in the paper. It would also be interesting to examine the role of other realistic features, such as the novice’s learning-by-doing and information asymmetry between different parties. We leave them as promising directions for future research.
@@ A.1 Apprenticeships Within a Firm
Suppose a firm (principal) interacts with two workers, and we call them expert ($i=1$) and novice ($i=2$). There is a divisible task to split between the two workers. The expert always has a knowledge level of $1$, and the novice starts with a knowledge level of 00.
Within a period, let $\alpha_{1}$ and $\alpha_{2}$ be the share of the task assigned to the expert and the novice, respectively ($\alpha_{1}+\alpha_{2}\leq 1$). If $\alpha_{i}$ share of the task is successful, it generates a revenue of $K\alpha_{i}$ (00 if unsuccessful). Each worker decides whether to exert an effort $e_{i}\in\{0,1\}$, unobservable to the firm, with cost $e_{i}c(\alpha_{i})$ for some strictly increasing and convex $c(\cdot)$. For each worker, given his knowledge level $s$ and effort $e$, the probability of success is $g(s)[eq+(1-e)p]$ with $01\geq q>p>$ 0, for some strictly increasing $g(\cdot)$.
@@
Let us further assume $K(q-p)=1$ and $g(s)=\frac{1+s}{2}$. In this case, the solution is always interior, with $\alpha_{1}^{*}(s)=\frac{3-s}{4}$ and $\alpha_{2}^{*}(s)=\frac{1+s}{4}$. Substituting them back into the payoff functions, we have
@@ A.2 Technology Transfers Between Competing Firms
Consider a government (principal) that wants to bring advanced technologies from foreign companies to its domestic market. Let Firm $1$ be an international firm with full knowledge and Firm 2 be a local firm with an initial knowledge level of 00.
@@ A.2.1 Homogeneous Products with a Cost Shifter
Suppose that the two firms produce the same product and engage in Cournot competition. The marginal cost of Firm 1 is zero, and the marginal cost of Firm 2 is $1-s$. Within a period, both firms simultaneously choose their quantities, denoted by $q_{1}$ and $q_{2}$. The inverse demand curve is $p=A-q_{1}-q_{2}$.
@@ B.1 Proofs for Sections 2 and 3
It is without loss to bound the principal’s payment from above by $\bar{p}\equiv\frac{w(0)-w(1)}{1-\delta}$, as it would have compensated the expert for his largest possible payoff losses at once. With this additional restriction, the domain of program (1) is a closed subset of a compact space ($[0,1]^{\infty}\times[0,\bar{p}]^{\infty}$, with the product topology), and therefore it is compact. Existence then follows from the extreme value theorem. ∎
@@
We prove that $(\bm{s},\ \bm{p}^{*}(\bm{s}))$ is implementable and weakly improves the principal’s time-00 profit, and that (P-IC) is slack in period $k+1$ under $(\bm{s},\ \bm{p}^{*}(\bm{s}))$.
Feasibility. Since $(\bm{s},\bm{p})$ is feasible, $(\bm{s},\ \bm{p}^{*}(\bm{s}))$ satisfies (M) and (LL), so it is feasible too.
@@
where the first inequality is (P-IC) at $t$ and the second inequality is (E-IC) at $t-1$.
@@
We haven shown that all (E-IC) constraints are binding under $(\bm{s},\ \bm{p}^{*}(\bm{s}))$, implying that the RHS above equals $\sum_{\tau=k+1}^{\infty}\delta^{\tau-(k+1)}p_{\tau}^{*}$. Therefore, (P-IC) is slack in period $k+1$ under $(\bm{s},\ \bm{p}^{*}(\bm{s}))$.
Principal’s profit is weakly improved. This is because the discounted bilateral surplus is solely determined by the knowledge sequence and is thus unchanged, but the payment scheme (P*) minimizes the expert’s time-00 value to $\frac{w(s_{0})}{1-\delta}$.
@@
We now perturb $\bm{s}$ to construct an implementable $(\bm{s}^{\prime},\ \bm{p}^{*}(\bm{s}^{\prime}))$ which strictly increases the principal’s time-00 profit.
@@
Constructing $\bm{s}^{\prime}$. Let $T\geq k+2$ be the first period after $k+1$ that has a higher knowledge level. That is, $s_{k+1}=\cdots=s_{T-1}<s_{T}$. Consider $\bm{s}^{\prime}$ such that $s_{t}^{\prime}=s_{t}+\varepsilon$ for $t\in\{k+1,\cdots,T-1\}$ and $s_{t}^{\prime}=s_{t}$ otherwise. Because (P-IC) is slack at $k+1$ under $(\bm{s},\ \bm{p}^{*}(\bm{s}))$, by continuity, it still holds under $(\bm{s}^{\prime},\ \bm{p}^{*}(\bm{s}^{\prime}))$ for a small positive $\varepsilon$. In particular, we can set $\varepsilon$ to be the maximum value in $[0,s_{T}-s_{T-1}]$ satisfying
@@
$(\bm{s}^{\prime},\ \bm{p}^{*}(\bm{s}^{\prime}))$ improves the principal’s profit. Since $G(\cdot)$ is strictly increasing, we have $\sum_{\tau=0}^{\infty}\delta^{\tau-t}G(s^{\prime}_{\tau})>\sum_{\tau=0}^{\infty}\delta^{\tau-t}G(s_{\tau})$, that is, the discounted bilateral surplus becomes strictly higher. Because the payment scheme (P*) anchors the expert’s time-00 value at $\frac{w(s_{0})}{1-\delta}$, the principal’s is strictly better off than under the original contract $(\bm{s},\bm{p})$.
@@
[…]
@@ B.2 Proofs for Section 4
Consider an overlapping generation model with a long-lived principal and a sequence of workers. At time 00, there are two workers, an expert and a novice, denoted by $i=-1$ and $i=0$, respectively. Fix an integer $K\geq 2$. At time $iK$, worker $i\geq 1$ arrives as a novice with a knowledge level of 00; worker $i-2$, who was the expert, exits the game; and worker $i-1$, who was the novice, becomes the expert. [^26] In other words, every worker $i\geq 0$ lives for $2K$ periods, with the first (last) $K$ periods being his novice (expert) phase; the initial expert, worker $i=-1$, only lives in the first $K$ periods of the game. In each period $t\in\{iK,iK+1,\cdots,(i+1)K-1\}$, the principal chooses a payment ${p}_{t}$ and the expert (i.e., worker $i-1$) chooses a training amount ${x}_{t}$, as illustrated in Figure 1; the novice (i.e., worker $i$) has no actions to take.
@@
which maps all the training amounts from period $iK$ to $t-1$ and the payments from period $(i+1)K$ to time $t-1$ to a current training amount. [^27] To be sure, each expert knows the training he received when he was a novice but did not observe the payments between the principal and the previous expert; therefore, $X_{iK}\times\cdots\times X_{(i+1)K-1}$ is in expert $i$’s information set, but $P_{iK}\times\cdots\times P_{(i+1)K-1}$ is not.
@@
[…]
@@
Let $\bm{s}^{\prime}$ be such that $s_{t}^{\prime}=s_{t}+\varepsilon$ for $t\in\{k,\cdots,T-1\}$ and $s_{t}^{\prime}=s_{t}$ otherwise. We verify that $(\bm{s}^{\prime},\ \bm{p}^{*}(\bm{s}^{\prime}))$ is implementable and strictly improves the principal’s profit, which will contradict the optimality of $(\bm{s},\bm{p})$. Profit is strictly improved because $\bm{s}^{\prime}$ dominates $\bm{s}$ and the expert’s time-00 value is fixed by (R-P*). For implementability, because (R-P*) binds all the (R-E-IC) constraints, it suffices to verify (R-S-IC) under $(\bm{s}^{\prime},\ \bm{p}^{*}(\bm{s}^{\prime}))$. The RHS of (R-S-IC) is unchanged for $t\in[1,k-1]\cup[T+1,K-1]$ and the LHS gets weakly higher, so these (R-S-IC) constraints are satisfied. In period $T$, (R-S-IC) holds because
@@
For part i), note that if $s_{1}^{*}=0$, then (R-BE) implies a sequence of zeros; if $s_{1}^{*}>0$, then (R-BE) implies a strictly increasing sequence. By Lemma 8, $s_{K}^{*}=1$, so we must have $10<s_{1}^{*}<\cdots<s_{K}^{*}=$ 1.

### v2

@@ 6.2 Direct Training Cost
6.3 Alternative Timing
In the baseline model, we assume delayed observation of the expert’s training, capturing that the novice’s knowledge growth can be difficult to verify instantly. Meanwhile, the principal’s payment in a given period is made and observed before the expert makes his decision. In this subsection, we consider an alternative timing under which the principal pays the expert after the expert chooses $x_{t}$, while we still maintain the assumption that $x_{t}$ is not observed until $t+1$. The new timing is shown in Figure 4.
$s_{t}$ is observedExpert gets $w(s_{t})+p_{t}$Principal gets $\pi(s_{t})-p_{t}$Expert chooses $x_{t}$Principal chooses $p_{t}$$s_{t+1}$ is observed($x_{t}$ is observed)
Figure 4: Alternative Timing of the Game.
Note that this timing is strategically equivalent to having the principal and expert act simultaneously within a period, because neither party can condition its move on the other party’s action within the same period. This implies that any consequences to either player’s deviation can only occur starting the next period. We have the following characterization of the optimal contract.
Theorem 4.
Let $(\bm{s}^{*},\ \bm{p}^{*})$ be the optimal contract obtained in Theorem 1, with discount factor $\delta^{2}$. Under the alternative timing, the unique optimal contract $\{(\tilde{x}_{t},\ \tilde{p}_{t})\}_{t\geq 0}$ has the following structure: For each $k\in\mathbb{N}$,
  1. i)
    $\tilde{p}_{2k}=0$ and $\tilde{x}_{2k}=s_{k+1}^{*}-s_{k}^{*}$ in every even period;
  2. ii)
    $\tilde{p}_{2k+1}=(1+\delta)p_{k+1}^{*}$ and $\tilde{x}_{2k+1}=0$ in every odd period.
Interestingly, the principal and the expert now have to make alternate moves in order to support the scheme. In every even period (including $t=0$), the expert transfers some knowledge to the novice, but the principal makes no payment. In every odd period, the principal compensates the expert for the present value of his payoff losses due to the knowledge transfer in the period before, but the expert will not provide any further training until the next period. [^23] The term $(1+\delta)$ appears because $\bm{p}^{*}$ is obtained using discount factor $\delta^{2}$ but $\tilde{p}_{2k+1}=\frac{w(s^{*}_{k})-w(s^{*}_{k+1})}{1-\delta}=(1+\delta)p^{*}_{k+1}$.
To understand this result, note that the optimal payment structure follows from the same reasoning as in the baseline model: Because $p_{0}$ does not affect any of the expert’s incentive constraints, it is optimally set to 0; moreover, the principal always wants to frontload payment and compensate the expert immediately after verifying each knowledge transfer. On the other hand, think from the expert’s perspective in period $1$. Because $x_{1}$ cannot be conditioned on $p_{1}$ (see Figure 4), $x_{1}$ can only provide incentives for the prior payment $p_{0}$, not for the current payment $p_{1}$. But since $\tilde{p}_{0}=0$, no such incentives are needed in period $1$, so any positive $x_{1}$ can be shifted to $x_{0}$ to accelerate knowledge transfer. That $\tilde{x}_{1}=0$ further implies that $\tilde{p}_{2}=0$, thereby kick-starting the alternating sequence of moves.
While the optimal contract shares similar qualitative features with the baseline model, the knowledge transfer process is slowed down by a half due to the alternating moves. Furthermore, because the effective discount factor for constructing the optimal contract is now $\delta^{2}$, the long-run transfer limit also becomes lower (Proposition 1). This suggests that the players’ sequence of moves in the stage game and how quickly the principal’s payment is verified can have a significant effect on the speed and eventual level of knowledge transfer.
@@ 7 Concluding Remarks
We study how a principal should incentivize an expert to share specialized knowledge with a novice, when the expert is concerned about the future competition he has to face from a more capable novice. In the optimal relational contract between the principal and the expert, the expert is asked to transfer as much knowledge as possible to the novice, without any spot payment, in the initial period. The lack of commitment on both sides necessitates gradual and perpetual knowledge transfers later on, with the principal providing lump-sum compensations to the expert for his future losses that arise from each transfer. When players are somewhat impatient, a complete knowledge transfer is not achievable even in the long run.
Our characterization matches well with the implementation details of the quid pro quo policy in the Chinese auto industry, as well as the development of Korea’s high-speed rail. In each of these cases, the government successfully facilitated significant technology transfers from various international companies to domestic manufacturers. Our analysis can also be extended to a more complex environment that accounts for expert retirements and the career progression of novices.
While the model is stylized, we believe that the insights can shed light on a variety of applications, within and beyond the examples discussed in the paper. It would also be interesting to examine the role of other realistic features, such as the novice’s learning-by-doing and the information asymmetry between different parties. We leave them as promising directions for future research.
Appendix
@@ A.1 Apprenticeships Within a Firm
Consider a firm (principal) interacting with two workers, and we call them expert ($i=1$) and novice ($i=2$). There is a divisible task to split between the two workers. The expert always has a knowledge level of $1$, and the novice starts with a knowledge level of 0.
Within a period, let $\alpha_{1}$ and $\alpha_{2}$ be the share of the task assigned to the expert and the novice, respectively ($\alpha_{1}+\alpha_{2}\leq 1$). If $\alpha_{i}$ share of the task is successful, it generates a revenue of $K\alpha_{i}$ (0 if unsuccessful). Each worker decides whether to exert an effort $e_{i}\in\{0,1\}$, unobservable to the firm, with cost $e_{i}c(\alpha_{i})$ for some strictly increasing and convex $c(\cdot)$. For each worker, given his knowledge level $s$ and effort $e$, the probability of success is g(s)[eq+(1e)p]g(s)[eq+(1-e)p] with $1\geq q>p>0$, for some strictly increasing $g(\cdot)$.
@@
Let us further assume K(qp)=1K(q-p)=1 and $g(s)=\frac{1+s}{2}$. In this case, the solution is always interior, with $\alpha_{1}^{*}(s)=\frac{3-s}{4}$ and $\alpha_{2}^{*}(s)=\frac{1+s}{4}$. Substituting them back into the payoff functions, we have
@@ A.2 Technology Transfers Between Competing Firms
Consider a government (principal) that wants to bring advanced technologies from foreign companies to its domestic market. Let Firm $1$ be an international firm with full knowledge and Firm 2 be a local firm with an initial knowledge level of 0.
@@ A.2.1 Homogeneous Products with a Cost Shifter
Suppose that the two firms produce the same product and engage in Cournot competition. The marginal cost of Firm 1 is zero, and the marginal cost of Firm 2 is 1s1-s. Within a period, both firms simultaneously choose their quantities, denoted by $q_{1}$ and $q_{2}$. The inverse demand curve is $p=A-q_{1}-q_{2}$.
@@ B.1 Proofs for Sections 2 and 3
It is without loss to bound the principal’s payment from above by $\bar{p}\equiv\frac{w(0)-w(1)}{1-\delta}$, as it would have compensated the expert for his largest possible payoff losses at once. With this additional restriction, the domain of program (1) is a closed subset of a compact space ($0,1]^{\infty}\times[0,\bar{p}]^{\infty}$, with the product topology), and therefore it is compact. Existence then follows from the extreme value theorem. ∎
@@
We prove that $(\bm{s},\ \bm{p}^{*}(\bm{s}))$ is implementable and weakly improves the principal’s time-0 profit, and that (P-IC) is slack in period $k+1$ under $(\bm{s},\ \bm{p}^{*}(\bm{s}))$.
Feasibility. Since $(\bm{s},\bm{p})$ is feasible, $(\bm{s},\ \bm{p}^{*}(\bm{s}))$ satisfies (M) and ($L$), so it is feasible too.
@@
where the first inequality is (P-IC) at $t$ and the second inequality is (E-IC) at t1t-1.
@@
We have shown that all (E-IC) constraints are binding under $(\bm{s},\ \bm{p}^{*}(\bm{s}))$, implying that the RHS above equals $\sum_{\tau=k+1}^{\infty}\delta^{\tau-(k+1)}p_{\tau}^{*}$. Therefore, (P-IC) is slack in period $k+1$ under $(\bm{s},\ \bm{p}^{*}(\bm{s}))$.
Principal’s profit is weakly improved. This is because the discounted bilateral surplus is solely determined by the knowledge sequence and is thus unchanged, but the payment scheme (P*) minimizes the expert’s time-0 value to $\frac{w(s_{0})}{1-\delta}$.
@@
We now perturb $\bm{s}$ to construct an implementable $(\bm{s}^{\prime},\ \bm{p}^{*}(\bm{s}^{\prime}))$ which strictly increases the principal’s time-0 profit.
@@
Constructing $\bm{s}^{\prime}$. Let $T\geq k+2$ be the first period after $k+1$ that has a higher knowledge level. That is, $s_{k+1}=\cdots=s_{T-1}<s_{T}$. Consider $\bm{s}^{\prime}$ such that $s_{t}^{\prime}=s_{t}+\varepsilon$ for $t\in\{k+1,\cdots,T-1\}$ and $s_{t}^{\prime}=s_{t}$ otherwise. Because (P-IC) is slack at $k+1$ under $(\bm{s},\ \bm{p}^{*}(\bm{s}))$, by continuity, it still holds under $(\bm{s}^{\prime},\ \bm{p}^{*}(\bm{s}^{\prime}))$ for a small positive $\varepsilon$. In particular, we can set $\varepsilon$ to be the maximum value in $0,s_{T}-s_{T-1}]$ satisfying
@@
$(\bm{s}^{\prime},\ \bm{p}^{*}(\bm{s}^{\prime}))$ improves the principal’s profit. Since $G(\cdot)$ is strictly increasing, we have $\sum_{\tau=0}^{\infty}\delta^{\tau-t}G(s^{\prime}_{\tau})>\sum_{\tau=0}^{\infty}\delta^{\tau-t}G(s_{\tau})$, that is, the discounted bilateral surplus becomes strictly higher. Because the payment scheme (P*) anchors the expert’s time-0 value at $\frac{w(s_{0})}{1-\delta}$, the principal’s is strictly better off than under the original contract $(\bm{s},\bm{p})$.
@@
A simple induction tells us that $s_{t}^{*}<\bar{s}^{*}$ for every $t$, so $\lim_{t}s_{t}^{*}\leq\bar{s}^{*}$. Together we have $\lim_{t}s_{t}^{*}=\bar{s}^{*}$, as desired. Parts iii) and iv) of Theorem 1 then follow directly from Lemma 3. Since $s_{1}^{*}>s_{0}=0$, Part i) follows from a simple induction on (BE).
@@ B.2 Proofs for Section 5
Consider an overlapping generation model with a long-lived principal and a sequence of workers. At time 0, there are two workers, an expert and a novice, denoted by i=1i=-1 and $i=0$, respectively. Fix an integer $K\geq 2$. At time $iK$, worker $i\geq 1$ arrives as a novice with a knowledge level of 0; worker i2i-2, who was the expert, exits the game; and worker i1i-1, who was the novice, becomes the expert. [^29] In other words, every worker $i\geq 0$ lives for $2K$ periods, with the first (last) $K$ periods being his novice (expert) phase; the initial expert, worker i=1i=-1, only lives in the first $K$ periods of the game. In each period $t\in\{iK,iK+1,\cdots,(i+1)K-1\}$, the principal chooses a payment ${p}_{t}$ and the expert (i.e., worker i1i-1) chooses a training amount ${x}_{t}$, as illustrated in Figure 1; the novice (i.e., worker $i$) has no actions to take.
@@
which maps all the training amounts from period $iK$ to t1t-1 and the payments from period $(i+1)K$ to time t1t-1 to a current training amount. [^30] To be sure, each expert knows the training he received when he was a novice but did not observe the payments between the principal and the previous expert; therefore, $X_{iK}\times\cdots\times X_{(i+1)K-1}$ is in expert $i$’s information set, but $P_{iK}\times\cdots\times P_{(i+1)K-1}$ is not.
@@
Now we further perturb $(\bm{s},\ \bm{p}^{*}(\bm{s}))$ to construct a strict improvement for the principal. Suppose that (R-S-IC) is slack in period $k\in\{1,\cdots,K-1\}$ under $(\bm{s},\ \bm{p}^{*}(\bm{s}))$. Let $T\geq k+1$ be the first period after $k$ such that $s_{T}>s_{k}$ (i.e., $s_{T}>s_{T-1}=\cdots=s_{k}$). [^31] TT is well-defined because i) by Lemma 7 it is without loss to set $s_{K}=1$; and ii) $s_{t}<1$ for every $1\leq t\leq K-1$, for otherwise the principal has no incentives to compensate the expert for the training $x_{t-1}$. Pick any $\varepsilon\in(0,s_{T}-s_{k})$ such that
@@
Let $\bm{s}^{\prime}$ be such that $s_{t}^{\prime}=s_{t}+\varepsilon$ for $t\in\{k,\cdots,T-1\}$ and $s_{t}^{\prime}=s_{t}$ otherwise. We verify that $(\bm{s}^{\prime},\ \bm{p}^{*}(\bm{s}^{\prime}))$ is implementable and strictly improves the principal’s profit, which will contradict the optimality of $(\bm{s},\bm{p})$. Profit is strictly improved because $\bm{s}^{\prime}$ dominates $\bm{s}$ and the expert’s time-0 value is fixed by (R-P*). For implementability, because (R-P*) binds all the (R-E-IC) constraints, it suffices to verify (R-S-IC) under $(\bm{s}^{\prime},\ \bm{p}^{*}(\bm{s}^{\prime}))$. The RHS of (R-S-IC) is unchanged for $t\in[1,k-1]\cup[T+1,K-1]$ and the LHS gets weakly higher, so these (R-S-IC) constraints are satisfied. In period $T$, (R-S-IC) holds because
@@
For part i), note that if $s_{1}^{*}=0$, then (R-BE) implies a sequence of zeros; if $s_{1}^{*}>0$, then (R-BE) implies a strictly increasing sequence. By Lemma 8, $s_{K}^{*}=1$, so we must have $0<s_{1}^{*}<\cdots<s_{K}^{*}=1$.
@@
B.3 Proofs for Section 6
Proof of Theorem 4.
Under the alternative timing, the expert’s IC constraints remain unchanged as (E-IC): $W_{t+1}(\bm{s},\bm{p})\geq\frac{w(s_{t})}{1-\delta},\ \forall t\geq 0$. However, since the expert can only react to the principal’s action in the next period, the principal’s IC constraints become:
$\Pi_{t}(\bm{s},\bm{p})\geq\pi(s_{t})+\frac{\delta\pi(s_{t+1})}{1-\delta},\quad\forall t\geq 0.$ (P-IC’)
The principal’s problem can now be written as:
[…]
Following an argument similar to that in the baseline model, one can prove an analogue of Claim 2: In any optimal contract, the time-0 payment must be zero and all the incentive constraints (except (P-IC) at $t=0$) must be binding. This has three implications. First, like before, the optimal payment stream is pinned down by (P*). Second, we obtain a new break-even condition:
$\delta^{2}[\pi(s_{t+2})-\pi(s_{t+1})]=w(s_{t-1})-w(s_{t}),\quad\forall t\geq 1.$ (BE’)
Third, we can rewrite the principal’s profit as
$\Pi_{0}(\bm{s},\bm{p})=\pi(s_{0})+\delta\pi(s_{1})+\frac{\delta^{2}\pi(s_{2})}{1-\delta}.$
Hence, the principal’s problem is equivalent to:
[…]
Let $\tilde{\bm{s}}$ be a solution to (13) with discount factor $\delta$, and let $\bm{s}^{*}$ be the optimal knowledge sequence obtained in Theorem 1 with discount factor $\delta^{2}$. Take any sequence $\bm{s}$ satisfying (M) and (BE’), and let $\bar{s}=\lim_{t}s_{t}$. Adding up (BE’) across all $t\geq 1$, we have
$\delta^{2}[\pi(\bar{s})-\pi(s_{2})]=w(s_{0})-w(\bar{s}).$
This indicates that $\bar{s}$ and $s_{2}$ mutually pin down each other, and that the choice of $s_{1}$ does not affect $\bar{s}$. Because the objective function is increasing in both $s_{1}$ and $s_{2}$, it is optimal to set $\tilde{s}_{1}=\tilde{s}_{2}=s_{1}^{*}$, where $s_{1}^{*}$ is defined by condition (4) using discount factor $\delta^{2}$.
But then, (BE’) recursively implies that $\tilde{s}_{2k+2}=\tilde{s}_{2k+1}$ for every $k\in\mathbb{N}$, that is, $\tilde{x}_{2k+1}=0$ in every odd period. Using this fact, one can further turn (BE’) into:
$\delta^{2}\left[\pi(\tilde{s}_{2k+2})-\pi(\tilde{s}_{2k})\right]=w(\tilde{s}_{2k-2})-w(\tilde{s}_{2k}),\quad\forall k\geq 1.$ (14)
Recall that $\bm{s}^{*}$ satisfies (BE) with discount factor $\delta^{2}$. Comparing (14) with (BE), we conclude that in every even period, we must have $\tilde{s}_{2k}=s^{*}_{k}$. Finally, the optimal payment sequence $\tilde{\bm{p}}$ is immediately implied by (P*). ∎